OpenStudy (anonymous):

(3x-1)^2=25 solve by using the square root method someone please help

OpenStudy (anonymous):

the answer is false

jimthompson5910 (jim_thompson5910):

\[\Large (3x-1)^2=25\] \[\Large 3x-1=\pm\sqrt{25}\] \[\Large 3x-1=\sqrt{25} \ \text{or} \ 3x-1=-\sqrt{25}\] \[\Large 3x-1=5 \ \text{or} \ 3x-1=-5\] \[\Large 3x=5+1 \ \text{or} \ 3x=-5+1\] \[\Large 3x=6 \ \text{or} \ 3x=-4\] \[\Large x=\frac{6}{3} \ \text{or} \ x=-\frac{4}{3}\] \[\Large x=2 \ \text{or} \ x=-\frac{4}{3}\]

OpenStudy (anonymous):

thank you so much!

jimthompson5910 (jim_thompson5910):

yw